OpenStudy (anonymous):

Write an equation in slope-intercept form for an equation of a line perpendicular to y = 24x +2 and that intersects at point (0, -2).

OpenStudy (radar):

Looking at your line equation the slope for that line is the coefficient of x, in this case slope m=24. A line perpendicular will have a slope that is the negative reciprocal or m=-1/24 Using a slope of -1/24 and the given point you can now develop the equation. Is this enough guidance?

OpenStudy (radar):

@allirose, answer please.

OpenStudy (anonymous):

it would be if i knew what i was doing but i dont know how to do this kind of math smh

OpenStudy (anonymous):

i really just need the answer please lol

OpenStudy (radar):

There are some methods that you must remember or have nearby as a reference. In this case you must know how slope is determined. \[\Delta y \over \Delta x\] . Delta X and Delta Y simply mean the change in y divided by the change in x. We will use this definition to solve your problem.

OpenStudy (radar):

No, you need more than the answer, that is another part of the problem. You need knowledge, that is why you are in school. But I digress, back to the problem -1/24=(y-(-2))/x-0) This is the point given in your problem. Solving The bottom line in the drawing is your answer Good luck you may need it.